OpenStudy (anonymous):

Write the quantity using sums and differences of simpler logarithmic expressions. a. log10(square root (x + 6)(x + 7)) b. ln (squared root(x + 6)(x + 7) divided by (squared root(x − 6)(x − 7)

myininaya (myininaya):

\[\ln(\frac{\sqrt{(x+6)(x+7)}}{\sqrt{(x-6)(x-7)}})=\ln(\sqrt{(x+6)(x+7)})-\ln(\sqrt{(x-6)(x-7)})\] \[=\ln[(x+6)(x+7)]^\frac{1}{2}-\ln[(x-6)(x-7)]^\frac{1}{2}\] \[=\frac{1}{2}*\ln[(x+6)(x+7)]-\frac{1}{2}*\ln[(x-6)(x-7)]\] \[=\frac{1}{2}*[\ln(x+6)+\ln(x+7)]-\frac{1}{2}*[\ln(x-6)+\ln(x-7)]\]

myininaya (myininaya):

is b

OpenStudy (anonymous):

\[\log_{10}\sqrt{(x+6)(x+7)}=\log(((x+6)(x+7))^{1/2})\]\[=1/2*(\log((x+6)(x+7))\]\[=1/2*[\log(x+6)+\log(x+7)]\]